# Hybrid Unsupervised Learning Strategy for Monitoring Industrial Batch Processes

### v1

@@ 1 Introduction
Continuous monitoring of complex manufacturing processes involves the systematic surveillance, measurement, and analysis of various aspects of a production system to ensure efficient operation, maintain product quality, and adhere to safety standards. Anomaly detection is a crucial aspect of a comprehensive monitoring system and involves identifying atypical patterns, events, or behaviors in dataset that significantly deviate from norm or expected values. In the domain of manufacturing processes, multivariate datasets encompass correlated time-sequenced records of process values. Multivariate anomaly detection techniques can be grouped into two main categories: statistical and machine learning (ML) methods (Schmidl et al., 2022). While production processes are characterized by high system order with complex interactions between multiple process values, machine learning (ML) methods are preferred – they can learn and model complex, non-linear relationships effectively (Helwig et al., 2015). Supervised ML methods require labelled training data (e.g. normal or abnormal behavior), whereas unsupervised do not. Employing unsupervised ML methods, it is assumed that the recorded data reflects the normal behavior of the process, thus reducing the cost of preparing and labelling data. Clustering-based ML techniques group similar data points into clusters - anomaly is measured by how well a data point fits into the identified clusters. An additional layer of analysis is added by interpreting the clusters as characteristic phases of a process (e.g., start, steady, transient, end). This can be useful when monitoring batch processes in the pharmaceutical industry, for example, where the duration and sequence of process phases (phase trajectory) are critical in terms of product quality. Self-organizing maps (SOMs), introduced by Kohonen (2013), have been successfully employed for process monitoring in various applications. While the approaches presented in Vermasvuori et al. (2002); Li et al. (2021) focus on anomaly detection, the concepts presented in von Birgelen and Niggemann (2017); Frey (2008) extend the utility of SOMs to identify, visualize, and monitor process phases. However, SOMs have certain limitations, especially when dealing with unbalanced training datasets and highly correlated process variables. To overcome these issues, a hybrid unsupervised learning strategy (HULS) have been presented by Frey (2023). In order to evaluate the performance of the developed method, a comparative study based on a synthetic dataset was conducted. The experiments showed that the HULS concept is significantly superior, especially in the reliable identification of unknown process phases. In this study, we explore the use of the HULS procedure for monitoring a laboratory batch process. Section II outlines the fundamental methodologies employed, while Section III gives an overview of the HULS approach. Section IV presents the experimental results, comparing the performance of HULS with SOM method. Finally, the paper is concluded in section V.
@@ 2.1 Self-Organizing Maps
SOMs are capable of mapping a high-dimensional input dataset $\mathbf{X}$ onto a low-dimensional grid while preserving the topological structure of the data. Usually structured in two layers, the input layer receives the input data points from the dataset $\mathbf{X}=\{\mathbf{x}_{1},...,\mathbf{x}_{L}\}$. Each data point is represented as an vector with $\mathbf{x}_{i}\in\mathbb{R}^{d}\ \forall\ i$ where $d$ corresponds to the number of dataset features. The output layer holds the neurons, typically arranged in a 2D lattice of size ${M\times N}$ with $P_{SOM}=MN$ neurons at position $\mathbf{v}\in\mathbb{N}=(1,...,M)\times(1,...,N)$. Each neuron is associated with a dedicated weight vector $\mathbf{w}_{\mathbf{v}}\in\mathbb{R}^{d}\ \forall\ \mathbf{v}$. The competitive learning rule is a key aspect of the SOM training process where neurons compete with each other to become the so called Best Matching Unit (BMU). The BMU is the neuron whose weight vector $\mathbf{w}_{\mathbf{v}}$ is the most similar to the input data point $\mathbf{x}_{i}$, whereby similarity is typically measured using the Euclidean distance
$e_{i\mathbf{v}}=\left\|\mathbf{x}_{i}-\mathbf{w}_{\mathbf{v}}\right\|,$ (1)
with the BMU index vector $\mathbf{v}^{*}$ calculated by:
$\mathbf{v}^{*}_{i}=\operatorname*{argmin}_{\mathbf{v}}\left\|\mathbf{x}_{i}-\mathbf{w}_{\mathbf{v}}\right\|.$ (2)
In the $k$th training epoch the weights of the BMU and its neighbors are updated to become more similar to the input vector by:
@@
where $\alpha(k)$ is the learning rate and $h_{\mathbf{v^{*}}\mathbf{v}}(k)$ is the neighborhood. Both functions can be applied in several ways (Kohonen, 2013), however they should be decreasing with $k$. With $\alpha_{0}$ being the initial learning rate and $K$ the number of total learning epochs, $\alpha(k)$ is calculated by:
@@
$h_{\mathbf{v^{*}}\mathbf{v}}(k)=\exp\left(-\frac{\left\|\mathbf{v}^{*}-\mathbf{v}\right\|}{\sigma_{0}\sigma(k)}\right),\sigma(k)=\exp\left(-\frac{k}{K}\right).$ (5)
@@
$E_{Q}=\frac{1}{L}\sum_{i=1}^{L}\left\|\mathbf{x}_{i}-\mathbf{w}_{\mathbf{v^{*}}}\right\|,E_{T}=\frac{1}{L}\sum_{i=1}^{L}\delta_{i}.$ (6)
@@ 2.2 Unified-Distance Matrix and Watershed Transformation
The unified-distance matrix (UM), presented by Ultsch and Siemon (1990), is valuable tool for visualizing and interpreting SOMs, providing insights into the underlying structure and clusters within the data. The base idea is to add a dimension to the SOM lattice, which reflects the summed distances between the individual neuron and its surrounding neighbors. With $\mathbf{v}=(m,n)$ be the position of the individual neuron and $\mathbf{p}=(m-1,...m+1)\times(n-1,...,n+1)$ be the position of the neighbors in the lattice the distance $U_{\mathbf{v}}$ is calculated by:
@@ 2.3 Instantaneous Topological Map
SOMs suffer from severe difficulties when the training data is unbalanced or the features are highly correlated, which is typical in an industrial manufacturing process. There are several approaches to overcome these problems (Ruppert, 2004), one of which is the Instantaneous Topological Map (ITM) developed by Jockusch and Ritter (1999). An ITM is also structured in two layers, where the input layer receives the data points $\mathbf{x}_{i}\in\mathbb{R}^{d}\ \forall\ i$ from the dataset $\mathbf{X}=\{\mathbf{x}_{1},...,\mathbf{x}_{L}\}$. The output layer holds the neurons with an associated weight vector $\mathbf{s}_{r}\in\mathbb{R}^{d}\ \forall\ r$ with a total of $P_{ITM}$ neurons and a set of edges (neighboring neurons) $Q_{j}$. An ITM has no fixed number of neurons $P_{ITM}$ and the creation of neurons is based on the relation between BMU $j^{*}$ and the second BMU $j^{**}$, which is determined e.g. by the Euclidean distance:
$j^{*}_{i}=\operatorname*{argmin}_{\mathbf{j}}\left\|\mathbf{x}_{i}-\mathbf{s}_{j}\right\|,\>j^{**}_{i}=\operatorname*{argmin}_{\mathbf{j\neq j^{*}}}\left\|\mathbf{x}_{i}-\mathbf{s}_{j}\right\|.$ (8)
@@
This property of an ITM can also be interpreted in the context of resampling He and Garcia (2009), where the weight vectors of the neurons $\mathbf{s}$ reflect the resampled dataset $\mathbf{X}$. To assess the quality of the ITM mapping, the quantification error $E_{Q}$ given in (7) can be employed, by using the BMU weight vectors $\mathbf{s}_{j^{*}}$.
@@
Output: ITM, weight vectors $\mathbf{s}$
@@ 3 Hybrid Unsupervised Learning Strategy
To address these challenges, in Frey (2023) a a hybrid unsupervised learning approach is proposed. The HULS concept combines the capabilities of SOMs and ITM. While ITM is also a topology preserving neural network it can handle correlated training data efficiently and re-sample the training data in an intuitive manner (Barreto et al., 2003). The output of an ITM is a set of neurons with associated weight vectors $\mathbf{s}_{j}$ and associated edges $Q_{j}$. However, unlike SOMs, ITMs do not map the input dataset $\mathbf{X}$ onto a 2D lattice and thus lacks an effective and intuitive clustering mechanism similar to e.g. UM based techniques. This is a key aspect considering the challenge of identifying and monitoring unknown process phases in an industrial manufacturing process. Therefore the proposed HULS concept for monitoring complex industrial manufacturing processes is based on the following idea: In the first step, the ITM algorithm is applied on the training dataset $\mathbf{X}$. Based on the threshold parameter $\beta$ of the ITM procedure the neuron generation is controlled, the resulting ITM neuron weights $\mathbf{s}$ can be understood as the resampled dataset of $\mathbf{X}$. In the next step, this set of ITM weight vectors $\mathbf{s}$ is then used used as training dataset for the SOM procedure. Based on the trained SOM model the subsequent steps of UM transformation and WT segmentation can be performed. Algorithm 3 summarizes the proposed HULS learning strategy.
@@ 4.2 Data acquisition
Data acquisition is the critical process of collecting and preparing diverse datasets, that are essential for training models to accurately reflect real-world scenarios. The model training phase uses this data to train, while the subsequent validation step tests the model on a separate dataset to ensureits ability to generalize and perform effectively on new, unseen data. For the following experiments it is assumed, that the training and validation datasets reflects the normal behavior of the process. With respect to this assumption, a training and a validation dataset, visualized in Fig. 2 (a), have been recorded. The multivariate dataset consists of the five process values (P,F,L,D,Z). These values are recorded over six batches, with batches T1 through T4 used for training and batches N1 and N2 used for validation. The length of the batches varies from 200 to 220 samples per batch. Note that the process values shown are normalized and the time reference is anonymized due to confidentiality restrictions.
@@ 4.3 Model training and validation
As a first result, Table 1 provides a comparison of the key quality measures for the two algorithms. The parameters are calculated based on the training batches T1 to T4. Comparing the quantization error $E_{Q}$ gives an idea of how well the two models perform in covering the training dataset. The proposed HULS method significantly outperforms the standard SOM. This phenomenon becomes clear when comparing Fig. 2 (b) and (c), which shows the time course of the Euclidean distance $e_{i\mathbf{v^{*}}}$ between the BMU and the training data samples $\mathbf{x}_{i}$. The increased $e_{i\mathbf{v^{*}}}$ for the standard SOM is particularly evident when fast dynamic transitions in the process behavior occur. This phenomenon is typical in real-world applications: training data samples that reflect the fast transition phases are underrepresented, resulting in an unbalanced dataset. In such situations, the learning algorithm of the standard SOM fails to adjust the weight vectors sufficiently to cover and accurately map the training data, resulting in increased error $e_{i\mathbf{v^{*}}}$. Furthermore, this effect becomes even more pronounced when considering the validation batches N1 and N2. The standard SOM does not generalize and perform effectively on new, unseen data when trained on an unbalanced dataset. In comparison the HULS procedure can handle an unbalanced dataset more efficiently, thus confirming its practical applicability and robustness for industrial process monitoring.
@@ 4.4 Discovering unknown process phases
Holistic process monitoring, especially in the pharmaceutical industry, involves several essential elements such as identifying unknown process phases, tracking their progression, determining the duration of each phase, and detecting any anomalies. A prerequisite for robust process phase monitoring is the reliable clustering of the training dataset. As described in section 2, the use of SOMs with UM transformation and WT segmentation represents a robust approach for unsupervised clustering. As already discussed, the considered laboratory can be taken representative for a wide range of production processes: strong correlation among the process values and distinct process phases with fast transitions between them. However, the SOM algorithm exhibits considerable sensitivity to the inherent attributes of the training dataset, potentially leading to an increased topological error $E_{T}$, for example. As previously outlined, the $E_{T}$ measures the ability of the mapping algorithm to accurately model the topological dependencies within the dataset. As shown in Table 1, the proposed HULS procedure again significantly outperforms the conventional SOM model. Evidently, a high topological error $E_{T}$ may cause the UM transformation to cluster the data insufficiently. This effect can be verified by examining the UM transformation and the WT segmentation. For the following comparative evaluation, both the UM transformation and the WT segmentation algorithms were executed using a watershed margin of $\phi=10$’. The resulting UM with the identified clusters using the standard SOM algorithm is shown as a 3D and 2D plot in Fig. 3 (a). The standard SOM algorithm was able to identify 3 three distinct process phases. This result contradicts the installed process flow control, with five distinct process phases implemented. In this simple example of the laboratory process, these phases can be read visually: by examining the time courses of the valve position (Z) in relation to the pump speed (D) in Fig. 2 (a) , five dedicated combinations can be read for each batch. In contrast, as shown in Fig. 3 (b), the HULS concept successfully and reliably identifies the five process phases. When comparing the obtained results, the efficiency of the developed HULS concept becomes apparent. Even under challenging conditions, such as strong correlation between the process variables and the presence of an unbalanced learning dataset, the HULS method is able to reliable identify the process phases. Another aspect related to $E_{T}$ is the efficiency how the intrinsic structure of the data is mapped in the model. A direct comparison of the two UM visualizations in Fig. 3 shows that the HULS model is more compact, i.e. fewer neurons are needed to store the data structure.

### v2

@@ 1 Introduction
Continuous monitoring of complex manufacturing processes involves the systematic surveillance, measurement, and analysis of various aspects of a production system to ensure efficient operation, maintain product quality, and comply with safety standards. Anomaly detection is a critical aspect of a comprehensive monitoring system and involves the identification of atypical patterns, events, or behaviors in datasets that significantly deviate from the norm or expected values. In the domain of manufacturing processes, multivariate datasets encompass correlated time-sequenced records of process values. Multivariate anomaly detection techniques can be divided into two main categories: statistical and machine learning (ML) methods (Schmidl et al., 2022). While production processes are characterized by high system order with complex interactions between multiple process values, machine learning (ML) methods are preferred - they can effectively learn and model complex, non-linear relationships (Helwig et al., 2015). Supervised ML methods require labeled training data (e.g., normal or abnormal behavior), while unsupervised methods do not. Unsupervised ML methods assume that the recorded data reflects the normal behavior of the process, thus reducing the cost of data preparation and labeling. Clustering-based ML techniques group similar data points into clusters - anomaly is measured by how well a data point fits into the identified clusters. An additional layer of analysis is added by interpreting the clusters as characteristic phases of a process (e.g., start, steady, transient, end). This can be useful when monitoring batch processes in the pharmaceutical industry, for example, where the duration and sequence of process phases (phase trajectory) are critical in terms of product quality. Self-organizing maps (SOMs), introduced by Kohonen (2013), have been successfully employed for process monitoring in various applications. While the approaches presented in Vermasvuori et al. (2002); Li et al. (2021) focus on anomaly detection, the concepts presented in von Birgelen and Niggemann (2017); Frey (2008) extend the utility of SOMs to identify, visualize, and monitor process phases. However, SOMs have certain limitations, especially when dealing with unbalanced training datasets and highly correlated process variables. To overcome these problems, a hybrid unsupervised learning strategy (HULS) have been presented by Frey (2023). In order to evaluate the performance of the developed method, a comparative study based on a synthetic dataset was conducted. The experiments showed that the HULS concept is significantly superior, especially in the reliable identification of unknown process phases. In this study, we explore the use of the HULS procedure for monitoring a laboratory batch process. Section 2 outlines the basic methods used, while Section 3 gives an overview of the HULS approach. Section 4 presents the experimental results, comparing the performance of HULS with SOM method. Finally, the paper is concluded in Section 5.
@@ 2.1 Self-Organizing Maps
SOMs are capable of mapping a high-dimensional input dataset $\mathbf{X}$ onto a low-dimensional grid while preserving the topological structure of the data. Usually structured in two layers, the input layer receives the input data points from the dataset $\mathbf{X}=\{\mathbf{x}_{1},...,\mathbf{x}_{L}\}$. Each data point is represented as an vector with $\mathbf{x}_{i}\in\mathbb{R}^{d}\ \forall\ i$ where $d$ corresponds to the number of dataset features. The output layer is typically a 2D grid of size ${M\times N}$ with $P_{SOM}=MN$ neurons at position $\mathbf{v}\in\mathbb{N}=(1,...,M)\times(1,...,N)$. The Best Matching Unit (BMU) is the neuron with index $\mathbf{v}^{*}_{i}=\operatorname*{argmin}_{\mathbf{v}}\left\|\mathbf{x}_{i}-\mathbf{w}_{\mathbf{v}}\right\|$ whose weight vector $\mathbf{w}_{\mathbf{v}}\in\mathbb{R}^{d}\ \forall\ \mathbf{v}$ is the most similar to data point $\mathbf{x}_{i}$, with similarity typically measured by the Euclidean distance:
$e_{i\mathbf{v}}=\left\|\mathbf{x}_{i}-\mathbf{w}_{\mathbf{v}}\right\|.$ (1)
In the $k$-th step of $K$ training epochs the weights of the BMU and its neighbors are updated by:
@@
where $\alpha(k)$ is the learning rate and $h_{\mathbf{v^{*}}\mathbf{v}}(k)$ is the neighborhood function. Both functions can be applied in several ways (Kohonen, 2013), however they should be decreasing with $k$. With $\alpha_{0}$ being the initial learning rate and $K$ the number of total learning epochs, $\alpha(k)$ is calculated by:
@@
$h_{\mathbf{v^{*}}\mathbf{v}}(k)=\exp\left(-\frac{\left\|\mathbf{v}^{*}-\mathbf{v}\right\|}{\sigma_{0}\sigma(k)}\right),\>\>\sigma(k)=\exp\left(-\frac{k}{K}\right).$ (4)
@@
$E_{Q}=\frac{1}{L}\sum_{i=1}^{L}\left\|\mathbf{x}_{i}-\mathbf{w}_{\mathbf{v^{*}}}\right\|$ (5)
$,\>\text{and}\>\>E_{T}=\frac{1}{L}\sum_{i=1}^{L}\delta_{i}.$ (6)
@@ 2.2 Unified-Distance Matrix and Watershed Transformation
The unified-distance matrix (UM), presented by Ultsch (2003), is valuable tool for visualizing and interpreting SOMs, providing insights into the underlying structure and clusters within the data. The base idea is to add a dimension to the SOM lattice, which reflects the summed distances between the individual neuron and its surrounding neighbors. With $\mathbf{v}=(m,n)$ be the position of the individual neuron and $\mathbf{p}=(m-1,...m+1)\times(n-1,...,n+1)$ be the position of the neighbors in the lattice the distance $U_{\mathbf{v}}$ is calculated by:
@@ 2.3 Instantaneous Topological Map
SOMs suffer from severe difficulties when the training data is unbalanced or when the features are highly correlated, which is typical in industrial manufacturing processes. There are several approaches to overcome these problems (Ruppert, 2004), one of which is the Instantaneous Topological Map (ITM) developed by Jockusch and Ritter (1999). An ITM is also structured in two layers, where the input layer receives the data points $\mathbf{x}_{i}\in\mathbb{R}^{d}\ \forall\ i$ from the dataset $\mathbf{X}=\{\mathbf{x}_{1},...,\mathbf{x}_{L}\}$. The output layer holds the neurons with an associated weight vector $\mathbf{s}_{r}\in\mathbb{R}^{d}\ \forall\ r$ with a total of $P_{ITM}$ neurons and a set of edges (neighboring neurons) $Q_{j}$. An ITM has no fixed number of neurons $P_{ITM}$ and the creation of neurons is based on the relation between BMU $j^{*}$ and the second BMU $j^{**}$, which is determined e.g. by the Euclidean distance:
$j^{*}_{i}=\operatorname*{argmin}_{\mathbf{j}}\left\|\mathbf{x}_{i}-\mathbf{s}_{j}\right\|$ (8)
$,j^{**}_{i}=\operatorname*{argmin}_{\mathbf{j\neq j^{*}}}\left\|\mathbf{x}_{i}-\mathbf{s}_{j}\right\|.$ (9)
@@
This property of an ITM can also be interpreted in the context of resampling He and Garcia (2009), where the set of neuron weight vectors $\mathbf{S}$ reflect the resampled dataset $\mathbf{X}$. To assess the quality of the ITM mapping, the quantification error $E_{Q}$ given in (5) can be employed, by using the BMU weight vectors $\mathbf{s}_{j^{*}}$.
@@
Output: ITM, set of weight vectors $\mathbf{S}$
@@ 3 Hybrid Unsupervised Learning Strategy
To address these challenges, in Frey (2023) a hybrid unsupervised learning approach is proposed. The HULS concept combines the capabilities of SOMs and ITM. While ITM is also a topology preserving neural network it can handle correlated training data efficiently and re-sample the training data in an intuitive manner (Barreto et al., 2003). The output of an ITM is a set of neurons with associated weight vectors $\mathbf{s}_{j}$ and associated edges $Q_{j}$. However, unlike SOMs, ITMs do not map the input dataset $\mathbf{X}$ onto a 2D lattice and thus lacks an effective and intuitive clustering mechanism similar to e.g. UM based techniques. This is a key aspect considering the challenge of identifying and monitoring unknown process phases in an industrial manufacturing process. Therefore the proposed HULS concept for monitoring complex industrial manufacturing processes is based on the following idea: In the first step, the ITM algorithm is applied on the training dataset $\mathbf{X}$. Based on the threshold parameter $\beta$ of the ITM procedure the neuron generation is controlled, the resulting ITM neuron weights $\mathbf{s}$ can be understood as the resampled dataset of $\mathbf{X}$. In the next step, this set of ITM weight vectors $\mathbf{S}$ is then used as training dataset for the SOM procedure. Based on the trained SOM model the subsequent steps of UM transformation and WT segmentation can be performed. Algorithm 3 summarizes the proposed HULS learning strategy.
@@ 4.2 Data acquisition
Data acquisition is the critical process of collecting and preparing diverse datasets, that are essential for training models to accurately reflect real-world scenarios. The model training phase uses these data to train, while the subsequent validation step tests the model on a separate dataset to ensure its ability to generalize and perform effectively on new, unseen data. For the following experiments it is assumed, that the training and validation datasets reflect the normal behavior of the process. With respect to this assumption, a training and a validation dataset, visualized in Fig. 2 (a), have been recorded. The multivariate dataset consist of the five process values (P,F,L,D,Z). These values are recorded over six batches, with batches T1 through T4 used for training and batches N1 and N2 used for validation. The length of the batches varies from 200 to 220 samples per batch. Note that the process values shown are normalized and the time reference is anonymized due to confidentiality restrictions.
@@ 4.3 Model training and validation
As a first result, Table 1 provides a comparison of the key quality measures for the two algorithms. The parameters are calculated based on the training batches T1 to T4. Comparing the quantization error $E_{Q}$ gives an idea of how well the two models perform in covering the training dataset. The proposed HULS method significantly outperforms the standard SOM. This phenomenon becomes clear when comparing Fig. 2 (b) and (c), which show the time courses of the Euclidean distance $e_{i\mathbf{v^{*}}}$ between the BMU and the training data samples $\mathbf{x}_{i}$. The increased $e_{i\mathbf{v^{*}}}$ for the standard SOM is particularly evident when fast dynamic transitions in the process behavior occur. This phenomenon is typical in real-world applications: training data samples that reflect the fast transition phases are underrepresented, resulting in an unbalanced dataset. In such situations, the learning algorithm of the standard SOM fails to adjust the weight vectors sufficiently to cover and accurately map the training data, resulting in increased error $e_{i\mathbf{v^{*}}}$. Furthermore, this effect becomes even more pronounced when considering the validation batches N1 and N2. The standard SOM does not generalize and perform effectively on new, unseen data when trained on an unbalanced dataset. In comparison the HULS procedure can handle an unbalanced dataset more efficiently, thus confirming its practical applicability and robustness for industrial process monitoring.
@@ 4.4 Discovering unknown process phases
Holistic process monitoring, especially in the pharmaceutical industry, involves several essential elements such as identifying unknown process phases, tracking their progression, determining the duration of each phase, and detecting any anomalies. A prerequisite for robust process phase monitoring is the reliable clustering of the training dataset. As described in section 2, the use of SOMs with UM transformation and WT segmentation represents a robust approach for unsupervised clustering. As already discussed, the considered laboratory can be taken representative for a wide range of production processes: strong correlation among the process values and distinct process phases with fast transitions between them. However, the SOM algorithm exhibits considerable sensitivity to the inherent attributes of the training dataset, potentially leading to an increased topological error $E_{T}$, for example. As previously outlined, the $E_{T}$ measures the ability of the mapping algorithm to accurately model the topological dependencies within the dataset. As shown in Table 1, the proposed HULS procedure again significantly outperforms the conventional SOM model. Evidently, a high topological error $E_{T}$ may cause the UM transformation to cluster the data insufficiently. This effect can be verified by examining the UM transformation and the WT segmentation.
For the following comparative evaluation, both the UM transformation and the WT segmentation algorithms were executed using a watershed margin of $\phi=10$’. The resulting UM with the identified clusters using the standard SOM algorithm is shown as a 3D and 2D plot in Fig. 3 (a). The standard SOM algorithm was able to identify three distinct process phases. This result contradicts the installed process flow control, with five distinct process phases implemented. In this simple example of the laboratory process, these phases can be read visually: by examining the time courses of the valve position (Z) in relation to the pump speed (D) in Fig. 2 (a), five dedicated combinations can be read for each batch. In contrast, as shown in Fig. 3 (b), the HULS concept successfully and reliably identifies the five process phases. When comparing the obtained results, the efficiency of the developed HULS concept becomes apparent. Even under challenging conditions, such as strong correlation between the process variables and the presence of an unbalanced learning dataset, the HULS method is able to reliable identify the process phases. Another aspect related to $E_{T}$ is the efficiency how the intrinsic structure of the data is mapped in the model. A direct comparison of the two UM visualizations in Fig. 3 shows that the HULS model is more compact, i.e. fewer neurons are needed to store the data structure.